### v1

# Many-hypercube codes: High-rate quantum error-correcting codes for high-performance fault-tolerant quantum computation

Abstract
Conventional approaches to quantum error correction for fault-tolerant quantum computation are based on encoding a single logical qubit into many physical qubits, resulting in asymptotically zero encoding rates and therefore huge resource overheads. To overcome this issue, high-rate quantum codes, such as quantum low-density parity-check codes, have been studied over the past decade. However, such codes have complex structure, making it difficult to perform logical gate operations in parallel without sacrificing their advantage. Observing the simple structure and high rates of quantum error-detecting codes, here we propose concatenated high-rate quantum error-detecting codes as a new family of high-rate quantum codes. Their simple structure allows for a geometrical interpretation using hypercubes, each of which corresponds to a logical qubit. We thus call them many-hypercube codes. The encoding rate is remarkably high, e.g., 30% (64 logical qubits are encoded into 216 physical qubits). Developing a dedicated high-performance decoder, we achieve high error thresholds even in a circuit-level noise model. Logical gate operations are also parallelizable. Thus, the many-hypercube codes will pave the way to high-performance fault-tolerant quantum computation.
I Introduction
Quantum computers have been expected to outperform current classical computers by harnessing quantum superposition states. However, the quantum superpositions are notoriously fragile. The so-called decoherence leads to many errors in quantum computers, spoiling quantum computation. A standard approach to this issue is quantum error correction. Careful use of quantum error-correcting codes can protect quantum computation from errors, which is called fault-tolerant quantum computation [1, 2, 3]. A standard quantum error-correcting code for this purpose is the surface code [4, 5, 6, 7, 8, 9, 10], which requires only a two-dimensional qubit array with nearest-neighbor interactions, and therefore is suitable for, e.g., superconducting-circuit implementations [11, 12, 13]. However, the surface code uses many physical qubits to protect a single logical qubit. More precisely, the number of physical qubits to encode a single logical qubit is the square of the code distance, where the code distance characterizes the code size and we can, in principle, correct arbitrary independent physical-qubit errors if the number of them is less than half the code distance. This means that the encoding rate defined as the ratio of the number of logical qubits to that of physical qubits vanishes rapidly as the code size becomes larger, resulting in large resource overheads [14, 15, 16].
In contrast, quantum low-density parity-check (qLDPC) codes [17] are known for their ability to achieve constant encoding rates, and therefore have been studied over the past decade to mitigate the above overhead issue [18, 19]. Various kinds of qLDPC codes have been proposed [20, 21, 22, 23, 24, 25, 26, 27, 28], high-performance decoders for them have also been developed [29, 30, 31, 32, 33, 34, 35], and their physical implementations have recently been proposed [36, 37, 38, 39]. However, they have relatively complex structure, making it difficult to implement logical gate operations in a fault-tolerant manner. A few methods for this purpose have been proposed [18, 37, 38, 39, 40, 41, 42], but parallel execution of logical gates maintaining the advantage of the qLDPC codes, i.e., low overheads is still challenging.
Recently, a completely different approach to the constant encoding rate has been proposed [43], which is based on a more conventional approach called code concatenation. Concatenation means recursive encoding with multiple codes. By increasing the encoding rates for higher concatenation levels, this proposal achieved a finite rate for the infinite code size. Importantly, this also allows for parallel execution of logical gates with constant overheads, unlike qLDPC codes. In other words, this proposal offers time-efficient, constant-space-overhead fault-tolerant quantum computation. However, this approach based on quantum Hamming codes has two issues. First, the encoding rate is not very high, converging to ${1/36}$. Second, the decoding of this concatenated codes is based on hard-decision decoding, which is known to be suboptimal and has relatively low performance.
In this work, we propose another family of high-rate concatenated quantum codes. The characteristic feature of our proposal is the use of quantum error-detecting codes [2], which have distance 2 and therefore can detect an error but cannot correct it. By concatenating the error-detecting codes, we can increase the code distance and thus obtain error-correcting codes. The advantage of the quantum error-detecting codes is their simple structure. Harnessing this advantage, Knill proposed concatenated quantum error-detecting codes called the ${C_{4}/C_{6}}$ scheme and achieved very high performance, i.e., the first realization of the error threshold exceeding 1% in a circuit-level noise model [44]. (Recently, the concatenation of the ${C_{4}/C_{6}}$ scheme and the concatenated quantum Hamming codes mentioned above has been proposed to improve the performance of the latter [45], but this still has the above-mentioned two issues, that is, the rate becomes rather lower and the decoding is still based on hard-decision decoding.) However, the ${C_{4}/C_{6}}$ scheme is based on a single-logical-qubit encoding, like the surface code, and therefore its encoding rate approaches zero rapidly as the code size increases. Unlike the ${C_{4}/C_{6}}$ scheme, we concatenate high-rate quantum error-detecting codes. In this work, we focus on the ${[\![6,4,2]\!]}$ code (see Appendix B for its details), where four qubits are encoded into six qubits and the distance is 2. (The reason for choosing this code is its relatively high rate and small size. The use of other codes, such as ${[\![4,2,2]\!]}$ and ${[\![8,6,2]\!]}$, is also interesting, but left for future work.) By concatenating it ${(L-1)}$ times, we obtain the ${[\![6^{L},4^{L},2^{L}]\!]}$ code, which we refer to as level-$L$ C642. Although it is not a constant-rate code, that is, its rate, ${(4/6)^{L}}$, approaches zero as $L$ becomes larger, the rate is remarkably high for small $L$, e.g., 30% (20%) at the level 3 (4) with distance 8 (16), which is higher than not only the surface code, but also well-studied qLDPC codes [36, 37, 38, 39]. Thus the C642, which we also call “many-hypercube codes” for the reason explained later (Fig. 1), will be promising as a near-term target. Note that the number of logical qubits is not limited to $K^{(L)}=4^{L}$. Using M code blocks encoded with the level-$L$ C642 as fault-tolerant quantum registers [43], we can use $MK^{(L)}$ logical qubits for fault-tolerant quantum computing.
We developed a high-performance decoder dedicated to the C642 based on level-by-level minimum distance decoding. (Our proposed decoding method will also be useful for other concatenated codes, such as the above concatenated quantum Hamming codes.) Using this decoder, we achieved a threshold of 5.6% for bit-flip errors, which is comparable to the surface code (10.9%) [46] and a 4%-rate qLDPC (hypergraph product) code (7.5%) [33]. We also propose fault-tolerant zero-state encoders for the C642. Using them, we achieved a threshold of 0.9% for a logical controlled-NOT (CNOT) gate in a circuit-level noise model. Finally, we explain how to perform logical gates for the C642 in parallel.
II Definition of C642 (many-hypercube codes)
Figure 1: Visualization of the structure of the level-3 C642 (many-hypercube code). The vertices correspond to physical qubits. Examples of logical $Z$ and $X$ operators and stabilizers are highlighted by blue ($Z$) and red ($X$). See the main text for the labelling. In general, logical $Z$ and $X$ operators for the level-$L$ C642 correspond to $L$-dimensional hypercubes, the number of which is equal to that of the logical qubits.
Using the ${[\![6,4,2]\!]}$ code recursively, e.g., the level-3 C642 is defined as follows. Each of physical and encoded qubits is labelled by three integers (corresponding to the level) and denoted by Q with subscripts of the integers. The corresponding operators are also labelled similarly. The four level-1 qubits $\{\mathrm{Q}_{i^{\prime},j^{\prime},k}^{(1)}|k=1,\ldots,4\}$ are encoded into the corresponding six physical qubits $\{\mathrm{Q}_{i^{\prime},j^{\prime},k^{\prime}}|k^{\prime}=1,\ldots,6\}$ ($i^{\prime},j^{\prime}=1,\ldots,6$), where the corresponding two stabilizers and logical $Z$ and $X$ operators are defined as follows (see Appendices A and B for the definitions of elementary gates and the ${[\![6,4,2]\!]}$ code, respectively):
$SZ_{i^{\prime},j^{\prime}}^{(1)}$ $=Z_{i^{\prime},j^{\prime},1}Z_{i^{\prime},j^{\prime},2}Z_{i^{\prime},j^{\prime},3}Z_{i^{\prime},j^{\prime},4}Z_{i^{\prime},j^{\prime},5}Z_{i^{\prime},j^{\prime},6},$ (1)
$SX_{i^{\prime},j^{\prime}}^{(1)}$ $=X_{i^{\prime},j^{\prime},1}X_{i^{\prime},j^{\prime},2}X_{i^{\prime},j^{\prime},3}X_{i^{\prime},j^{\prime},4}X_{i^{\prime},j^{\prime},5}X_{i^{\prime},j^{\prime},6},$ (2)
$Z_{i^{\prime},j^{\prime},1}^{(1)}$ $=Z_{i^{\prime},j^{\prime},1}Z_{i^{\prime},j^{\prime},2},~{}X_{i^{\prime},j^{\prime},1}^{(1)}=X_{i^{\prime},j^{\prime},2}X_{i^{\prime},j^{\prime},3},$ (3)
$Z_{i^{\prime},j^{\prime},2}^{(1)}$ $=Z_{i^{\prime},j^{\prime},2}Z_{i^{\prime},j^{\prime},3},~{}X_{i^{\prime},j^{\prime},2}^{(1)}=X_{i^{\prime},j^{\prime},1}X_{i^{\prime},j^{\prime},2},$ (4)
$Z_{i^{\prime},j^{\prime},3}^{(1)}$ $=Z_{i^{\prime},j^{\prime},4}Z_{i^{\prime},j^{\prime},5},~{}X_{i^{\prime},j^{\prime},3}^{(1)}=X_{i^{\prime},j^{\prime},5}X_{i^{\prime},j^{\prime},6},$ (5)
$Z_{i^{\prime},j^{\prime},4}^{(1)}$ $=Z_{i^{\prime},j^{\prime},5}Z_{i^{\prime},j^{\prime},6},~{}X_{i^{\prime},j^{\prime},4}^{(1)}=X_{i^{\prime},j^{\prime},4}X_{i^{\prime},j^{\prime},5}.$ (6)
Similarly, the four level-2 qubits $\{\mathrm{Q}_{i^{\prime},j,k}^{(2)}|j=1,\ldots,4\}$ are encoded into the corresponding six level-1 qubits $\{\mathrm{Q}_{i^{\prime},j^{\prime},k}^{(1)}|j^{\prime}=1,\ldots,6\}$ ($i^{\prime}=1,\ldots,6,k=1,\ldots,4$) and the four level-3 qubits $\{\mathrm{Q}_{i,j,k}^{(3)}|i=1,\ldots,4\}$ are encoded into the corresponding six level-2 qubits $\{\mathrm{Q}_{i^{\prime},j,k}^{(2)}|i^{\prime}=1,\ldots,6\}$ ($j,k=1,\ldots,4$). In this case, we use the 64 level-3 encoded qubits as logical qubits as a fault-tolerant quantum register for quantum computing. In general, the level-$L$ C642 is defined similarly.
Interestingly, the above code structure can be visualized as shown in Fig. 1, where the logical $Z$ and $X$ operators correspond to cubes. In general, the logical $Z$ and $X$ operators of the level-$L$ C642 correspond to $L$-dimensional hypercubes in similar visualization, where the number of hypercubes is equal to that of logical qubits. We thus call the C642 “many-hypercube codes.” [The “many” is used to distinguish it from a conventional hypercube code [47], which is the ${[\![2^{D},D,2]\!]}$ code defined on a single $D$-dimensional hypercube.] The number of the vertices of each $L$-dimensional hypercube is $2^{L}$, which is equal to the code distance. That is, the size of the many-hypercube codes is increased by increasing not the edge length but the dimension while keeping the edge length. In contrast, the logical operators and stabilizers of topological codes, such as the surface code [4, 5, 6, 7, 8, 9, 10] and the color code [48, 49, 50, 51], correspond to the edges and faces, respectively, and their code sizes are increased by increasing the edge length. Unlike the topological codes, the many-hypercube codes require interactions beyond nearest-neighbor ones. This is experimentally challenging, but recent experimental advances in ion-trap [52, 53, 54] and neutral-atom [55, 56, 57, 58] systems are encouraging. In fact, hypercube connectivity has already been realized experimentally [58].
III Decoding of C642
In this work, we investigated three decoding methods for the C642. The first one is Knill’s method proposed for the ${C_{4}/C_{6}}$ scheme [44], which we refer to as hard-decision decoding. The second one is soft-decision decoding based on a posteriori probability calculation [59, 60]. We modified it for the C642 because it is not directly applicable to high-rate codes. The third one is our proposed method based on level-by-level minimum distance decoding, where we keep only minimum-distance codewords from measurement outcomes as candidates at each level from level 1 to the logical level. See Appendices DF for the details of the three methods.
In the following, we assume that error correction is done by error-correcting teleportation (ECT) [44, 60, 61], where the decoding is the task to estimate the measurement outcomes of logical qubits from those of physical qubits in the logical Bell measurements (see Appendix C). Note that in ECT, we obtain all the syndrome information at once, no need to measure them one by one, unlike the surface code and qLDPC codes. We also assume that all the measurements are done in the $Z$ basis ${\{|0\rangle,|1\rangle\}}$, resulting in bit-value outcomes.
We numerically evaluated the performance of the three decoding methods using a bit-flip error model (see Appendix G for details). In this work, stabilizer simulations in our numerical study were done using a python package called Stim [62]. The results are shown in Fig. 2. First, the exponents of the power-function fits for the hard-decision decoding are smaller than half the code distance, as in the case of the ${C_{4}/C_{6}}$ scheme [60], which shows its suboptimality. On the other hand, the exponents for the other two are near to or even exceed half the distance, showing their high performance. Second, the error thresholds are estimated at 1.1%, 1.5%, and 5.6%, respectively, and only our minimum distance decoding is comparable to the surface code (10.9%) [46] and a well-studied qLDPC (4%-rate hypergraph product) code (7.5%) [33]. Thus our proposed decoding method achieves the highest performance among the three methods.
Figure 2: Performance of three decoders of the C642 for bit-flip errors. (a) Hard-decision decoder. (b) Symbol-MAP soft-decision decoder. (c) Level-by-level minimum distance decoder. The exponents are estimated by fitting a power function to the linear parts of the log-log plots. The thresholds indicated by arrows are defined by the intersection points of the level-3 and level-4 curves. See Appendices DF for the details of the three decoding methods and Appendix G for the simulation.
IV Fault-tolerant zero-state encoders of C642
Hereafter, we consider a circuit-level noise model with error rate $p_{\mathrm{circ}}$. In this work, we assume the following noise model. Each of physical-qubit zero-state preparation and Z-basis measurement is accompanied by a bit flip with probability $p_{\mathrm{circ}}$, and each two-qubit (CNOT) gate is followed by one of 15 two-qubit Pauli errors with equal probability $p_{\mathrm{circ}}/15$ [8]. On the other hand, we assume no single-qubit-gate and memory errors for simplicity. This model is relevant for ion-trap [52, 53, 54] and neutral-atom [55, 56, 57, 58] systems.
In such a circuit-level noise model, we cannot use the non-fault-tolerant encoders in Figs. 5d and 5e in Appendix B, because of uncorrectable correlated errors due to CNOT gates in a code block. Therefore we modify the zero-state encoder at each level as follows. In the following, we use various kinds of error detection and repeat encoding until no error is detected.
At level 1, we can use the well-known fault-tolerant GHZ-state preparation method with an ancilla qubit shown in Fig. 3a. We cannot apply this method directly to level 2, because of undetectable intra-block errors. To achieve fault tolerance with minimum effort at level 2, we propose the level-2 zero-state encoder shown in Fig. 3b, where level-1 error-detection gadgets are implemented by the flag-based method [63] (see Figs 12a and 12b in Appendix H). The transversal encoded CNOT and Hadamard gates can be performed fault-tolerantly by transversal physical gates, as shown in Figs. 5a and 5b in Appendix B. The level-3 zero-state encoder is the same as Fig. 3b if the levels are raised by one, but level-2 error-detection gadgets are implemented by the Steane method [64] (see Figs 12c and 12d in Appendix H). To improve the performance at level 4, we propose the level-4 zero-state encoder using two level-3 four-qubit GHZ states and error-detecting teleportation gadgets shown in Fig. 3c. As numerically shown below, these encoders well satisfy fault tolerance.
[…]
V Logical gate operations for C642
Arbitrary encoded Pauli gates can easily be performed fault-tolerantly by physical Pauli gates (or the so-called Pauli frame [44, 61]) according to the definitions of the encoded Pauli operators. For the other gates necessary for universal computation (see Appendix A), we first consider the case where the same gate is performed on all the encoded/physical qubits in a code block, which we refer to as full transversal gates. If the same gate is performed on only some specified qubits in a code block, we call it partial transversal gates, which is harder to implement.
From Figs. 5a and 5b in Appendix B, full transversal logical CNOT and Hadamard gates can be performed fault-tolerantly by full transversal physical CNOT and Hadamard gates (and physical SWAP gates) followed by ECT gadgets. We numerically evaluated the performance of the full transversal logical CNOT gates using the above noise model (see Appendix I for details). The results are shown in Fig. 4. The error threshold is estimated at 0.9%, which is comparable to the surface code (1.1%) [65] and a recently developed qLDPC quantum memory (0.8%) [39]. Also, the exponents of the power-function fits are close to half the code distance, showing that the logical CNOT gates (and the zero-state encoders in Fig. 3) well satisfy fault tolerance.
Figure 4: Performance of full transversal logical CNOT gates for the C642 in the circuit-level error model. Each circle shows the error probability per logical CNOT gate for error rate $p_{\mathrm{circ}}$. Error bars are negligible compared with the symbol size except for the level-4 result at $p_{\mathrm{circ}}=10^{-3}$. The dotted lines are power-function fits to the five data points from the lowest, from which the exponents are estimated. See Appendix I for the details of the simulation and the main text for the definition of the error model.
Next, we consider full transversal logical phase and non-Clifford gates, $S$ and ${R_{Y}(\pi/4)}$ (see Appendix A for their definitions). It is known that these two gates can be performed using ancilla qubits in $|Y\rangle=(|0\rangle+i|1\rangle)/\sqrt{2}$ and $|H\rangle=\cos(\pi/8)|0\rangle+\sin(\pi/8)|0\rangle$, respectively, together with CNOT and Hadamard gates, where $|Y\rangle$ and $|H\rangle$ are the eigenstates of $Y$ and $H$ with eigenvalue 1, respectively [15, 66] (see Fig. 14 in Appendix J). Therefore the full transversal logical $S$ and ${R_{Y}(\pi/4)}$ require to fault-tolerantly prepare the level-$L$ logical $|Y\rangle_{L}^{\otimes K^{(L)}}$ and $|H\rangle_{L}^{\otimes K^{(L)}}$, which can be achieved by the standard state distillation methods only with the arbitrary-state encoder in Fig. 5e in Appendix B and full transversal encoded CNOT and Hadamard gates [15, 66]. For the transversal logical non-Clifford gates, however, we also need partial transversal logical Hadamard gates depending on the measurement outcomes (see Fig. 14b in Appendix J).
The partial transversal logical Hadamard gates can be performed using an ancilla register where some logical qubits are in $|+\rangle=H|0\rangle=(|0\rangle+|1\rangle)/\sqrt{2}$ and the others in $|0\rangle$, which we refer to as a partial |+ket|$\rangle$| + ⟩ state. The partial |+ket|$\rangle$| + ⟩ state can be prepared fault-tolerantly (see Appendix K). From Figs. 14a and 14b in Appendix J, partial transversal logical phase and non-Clifford gates can also be performed using partial transversal logical Hadamard gates.
Partial transversal logical CNOT gates between two registers (inter-block logical CNOT gates) can be performed using two ancilla registers in the same partial |+ket|$\rangle$| + ⟩ state and full transversal logical CNOT and Hadamard gates (see Fig. 16a in Appendix M). In this case, each logical CNOT gate is performed on the corresponding two logical qubits with the same label in the two registers. To perform a logical CNOT gate on two logical qubits with different labels, we can use logical SWAP gates (see Appendix L), which allow us to perform a logical CNOT gate on an arbitrary logical-qubit pair between the two registers. Finally, logical CNOT gates in a register (intra-block logical CNOT gates) can be performed with an additional ancilla register in the logical all-zero state, logical SWAP gates, and full and partial transversal logical CNOT gates (see Fig. 16b in Appendix M).
Thus, we can implement a universal gate set (Pauli, Hadamard, phase, non-Clifford, and CNOT gates) for the C642 fault-tolerantly and perform most of them in parallel.
VI Summary and outlook
We have proposed concatenated high-rate quantum error-detecting codes as a new family of high-rate quantum codes for fault-tolerant quantum computation. Since the simple code structure allows for a geometrical interpretation using hypercubes corresponding to logical qubits, we call them many-hypercube codes. The encoding rates are remarkably high, 30% and 20% for the code distances of 8 and 16, respectively. We have developed a high-performance decoder and fault-tolerant zero-state encoders dedicated to the codes. Using them, we have achieved high error thresholds: 5.6% for bit-flip errors and 0.9% for a logical CNOT gate in a circuit-level noise model. Further improvements of the decoder and encoders for the many-hypercube codes are challenging but desirable. We have also explained how to implement logical gates necessary for universal quantum computation. More efficient logical-gate implementations for the many-hypercube codes may be possible but are left for future work. Minimizing the number of logical gates and computational depth for given quantum circuits is also an important compilation problem for the many-hypercube codes.
Acknowledgements.
Parts of numerical simulations were done on the HOKUSAI supercomputer at RIKEN (Project ID Q23614, RB230022).
Appendix A Elementary gates
In the $Z$ basis $\{|0\rangle,|1\rangle\}$, the matrix representations of elementary gates for universal quantum computation are as follows [3]:
Paul $\mbox{Paul gates}:X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},$ (11)
$\qquad\qquad\qquad Y=iXZ=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}$ (14)
Hadamard $\mbox{Hadamard gate}:H=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$ (17)
Phase $\mbox{Phase gate}:S=\begin{pmatrix}1&0\\ 0&i\end{pmatrix}$ (20)
[…]
CNOT $\mbox{CNOT gate}:U_{\mathrm{CNOT}}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix}$ (28)
Appendix B The ${[\![6,4,2]\!]}$ code
The ${[\![6,4,2]\!]}$ code is one of the simplest stabilizer codes [3] and has only two stabilizer generators (check operators): $S_{Z}=Z_{1}Z_{2}Z_{3}Z_{4}Z_{5}Z_{6}$ and $S_{X}=X_{1}X_{2}X_{3}X_{4}X_{5}X_{6}$, which can detect an $X$ (bit-flip) error and a $Z$ (phase-flip) error, respectively. Its four logical qubits are defined by the following logical $Z$ and $X$ operators:
$Z_{L1}$ $=Z_{1}Z_{2},~{}X_{L1}=X_{2}X_{3},$ (29)
$Z_{L2}$ $=Z_{2}Z_{3},~{}X_{L2}=X_{1}X_{2},$ (30)
$Z_{L3}$ $=Z_{4}Z_{5},~{}X_{L3}=X_{5}X_{6},$ (31)
$Z_{L4}$ $=Z_{5}Z_{6},~{}X_{L4}=X_{4}X_{5}.$ (32)
Other definitions are possible [2], but we use this because of the geometrical interpretation of the code structure using hypercubes (Fig. 1). From the definition, we can easily find that transversal logical CNOT and Hadamard gates can be performed by transversal physical CNOT and Hadamard (and SWAP) gates, respectively, as shown in Figs. 5a and 5b. The logical SWAP gates between logical qubits 1 and 2, 3 and 4, and (1,2) and (3,4) can also be performed by physical SWAP gates, as shown in Fig. 5c. The logical all-zero state of the ${[\![6,4,2]\!]}$ code is the six-qubit GHZ state, i.e.,
$|0000\rangle_{L}=\frac{|000000\rangle+|111111\rangle}{\sqrt{2}}.$ (33)
Thus the zero-state encoder of the ${[\![6,4,2]\!]}$ code is given by, e.g., the quantum circuit shown in Fig. 5d. This can be generalized to an arbitrary-state encoder by adding two CNOT gates, as shown in Fig. 5e.
Figure 5: Logical gates and encoders of the ${[\![6,4,2]\!]}$ code. (a) Transversal logical CNOT gates. (b) Transversal logical Hadamard gates. (c) Logical SWAP gates. In (a–c), the left and right sides correspond to logical and physical ones, respectively. (d) Zero-state encoder. (e) Arbitrary-state encoder. In (e), additional two CNOT gates are highlighted in red.
Appendix C Error-correcting teleportation
Our quantum error correction is based on error-correcting teleportation (ECT) [44, 60, 61]. This is quantum teleportation with logical qubits, where the classical feed-forward information is determined by decoding physical-qubit measurement outcomes, as shown in Fig. 6. Assuming that the two ancilla registers are reliable sufficiently, the errors come mainly from the failure of the decoding.
Figure 6: Error-correcting teleportation. The input logical-qubit state encoded with the level-$L$ C642, $|\psi\rangle_{L}$, on the left side is teleported with two logical all-zero states to the right side. The gray bold allows show the flow of the classical information. The “D” boxes are the decoders which decode the physical-qubit measurement outcomes and provide logical-qubit ones. If a logical-qubit outcome is 1, the logical $Z$ or $X$ operator is performed on the corresponding logical qubit. The logical CNOT and Hadamard gates are performed by transversal physical CNOT and Hadamard gates, as shown in Figs. 5a and 5b.
Appendix D Hard-decision decoding of C642
The hard-decision decoding of, e.g., the level-3 C642 is done as follows [44]. The measurement outcomes of physical qubits in each level-1 block and the corresponding level-1 outcomes are denoted by $\{x_{i^{\prime},j^{\prime},k^{\prime}}|k^{\prime}=1,\ldots,6\}$ and $\{x_{i^{\prime},j^{\prime},k}^{(1)}|k=1,\ldots,4\}$, respectively. All the four $x_{i^{\prime},j^{\prime},k}^{(1)}$ are set to $F$ (flag indicating an error) if the parity of the six bits, $x_{i^{\prime},j^{\prime},k^{\prime}}$, is odd according to the $Z$-stabilizer. Otherwise, they are set according to the definition of the logical $Z$ operators as follows: $x_{i^{\prime},j^{\prime},1}^{(1)}=x_{i^{\prime},j^{\prime},1}+x_{i^{\prime},j^{\prime},2}$, $x_{i^{\prime},j^{\prime},2}^{(1)}=x_{i^{\prime},j^{\prime},2}+x_{i^{\prime},j^{\prime},3}$, $x_{i^{\prime},j^{\prime},3}^{(1)}=x_{i^{\prime},j^{\prime},4}+x_{i^{\prime},j^{\prime},5}$, and $x_{i^{\prime},j^{\prime},4}^{(1)}=x_{i^{\prime},j^{\prime},5}+x_{i^{\prime},j^{\prime},6}$ (mod 2). Next, using the level-1 outcomes in each level-2 block, $\{x_{i^{\prime},j^{\prime},k}^{(1)}|j^{\prime}=1,\ldots,6\}$, we obtain the corresponding level-2 outcomes, $\{x_{i^{\prime},j,k}^{(2)}|j=1,\ldots,4\}$, as follows. If there is a single $F$ in the six $x_{i^{\prime},j^{\prime},k}^{(1)}$, e.g., $x_{i^{\prime},1,k}^{(1)}$, we can correct this as $x_{i^{\prime},1,k}^{(1)}=x_{i^{\prime},2,k}^{(1)}+x_{i^{\prime},3,k}^{(1)}+x_{i^{\prime},4,k}^{(1)}+x_{i^{\prime},5,k}^{(1)}+x_{i^{\prime},6,k}^{(1)}$ (mod 2) according to the $Z$-stabilizer. (Note that error-detecting codes can correct a located error.) Then, the four $x_{i^{\prime},j,k}^{(2)}$ are set according to the definition of the logical $Z$ operator, as above. If there is no $F$ and the parity of the six bits is even, we set the four $x_{i^{\prime},j,k}^{(2)}$ similarly. Otherwise, we set all the four $x_{i^{\prime},j,k}^{(2)}$ to $F$. Applying this decoding recursively, we finally obtain the logical-level outcomes. If it is $F$, then we randomly choose 0 or 1. For error detection used in the proposed encoders, the decoder returns $F$ (indicating detected errors) unless no $F$ is obtained throughout the decoding.
Appendix E Symbol-MAP soft-decision decoding of C642
The original proposal of the soft-decision decoding [60] is not directly applicable to the C642, because this decoding is based on block-MAP (maximum a posteriori probability) decoding, but the C642 has an exponentially large number of logical-qubit values in a code block and therefore we cannot evaluate all the corresponding a posteriori probabilities efficiently. So, we modify this to a symbol-MAP version, where we calculate marginal probabilities for each logical qubit, which can be done efficiently. The symbol-MAP version of the soft-decision decoding of, e.g., the level-3 C642 is as follows. Using the prior physical-qubit error probability $p_{e}$ and the physical-qubit measurement outcomes $\{x_{i^{\prime},j^{\prime},k^{\prime}}|k^{\prime}=1,\ldots,6\}$, the a posteriori probability for a physical-qubit value $x$ is expressed as
$p_{i^{\prime},j^{\prime},k^{\prime}}(x)=p_{e}^{I(x\neq x_{i^{\prime},j^{\prime},k^{\prime}})}(1-p_{e})^{I(x=x_{i^{\prime},j^{\prime},k^{\prime}})},$ (34)
where $I(A)$ is the indicator function that returns 1 if $A$ is true and otherwise 0. Then, the marginal a posteriori probability for a level-1 qubit, e.g., $\mathrm{Q}_{i^{\prime},j^{\prime},1}^{(1)}$, is given by
$p_{i^{\prime},j^{\prime},1}^{(1)}(x)=\frac{\displaystyle\sum_{x_{2}=0,1}\sum_{x_{3}=0,1}\sum_{x_{4}=0,1}R_{i^{\prime},j^{\prime}}^{(1)}(x,x_{2},x_{3},x_{4})}{\displaystyle\sum_{x_{1}=0,1}\sum_{x_{2}=0,1}\sum_{x_{3}=0,1}\sum_{x_{4}=0,1}R_{i^{\prime},j^{\prime}}^{(1)}(x_{1},x_{2},x_{3},x_{4})},$ (35)
$R_{i^{\prime},j^{\prime}}^{(1)}(x_{1},x_{2},x_{3},x_{4})$
$=\sum_{x^{\prime}_{1}=0,1}\sum_{x^{\prime}_{2}=0,1}\sum_{x^{\prime}_{3}=0,1}\sum_{x^{\prime}_{4}=0,1}\sum_{x^{\prime}_{5}=0,1}\sum_{x^{\prime}_{6}=0,1)}\prod_{k^{\prime}=1}^{6}p_{i^{\prime},j^{\prime},k^{\prime}}(x^{\prime}_{k^{\prime}})$
×$\qquad\times I\left(\sum_{k=1}^{6}x^{\prime}_{k}=0\right)I(x_{1}=x^{\prime}_{1}+x^{\prime}_{2})I(x_{2}=x^{\prime}_{2}+x^{\prime}_{3})$
×$\qquad\times I(x_{3}=x^{\prime}_{4}+x^{\prime}_{5})I(x_{4}=x^{\prime}_{5}+x^{\prime}_{6}),$ (36)
where the summation in the indicator functions is modulo 2. Repeating such calculations recursively, we finally obtain the marginal a posteriori probabilities for the logical-qubit values $\{x_{i,j,k}^{(3)}\}$. If $p_{i,j,k}^{(3)}(0)>0.5$, we set $x_{i,j,k}^{(3)}$ to 0 and otherwise to 1.
Appendix F Level-by-level minimum distance decoding of C642
The minimum distance decoding is known as one of the highest-performance decoding methods, where we search for a codeword closest to the measurement outcomes in the sense of the Hamming distance. However, there are an exponentially large number of codewords in general, hence it is usually intractable to find such a minimum-distance codeword. To solve this issue, we keep only minimum-distance codewords and discard larger-distance ones at each level from level 1 to the logical level. Since each level-1 codeword [a bit string corresponding to a level-1 computational-basis state, e.g., 000000 or 111111 for $|0000\rangle_{L}$, see Eq. (33)] consists of only six bits, we can easily select minimum-distance codewords and corresponding encoded bit strings (e.g., $0$ for $|0000\rangle_{L}$) in each level-1 block. While there is a single minimum-distance codeword (encoded bit string) when the parity of the six bits is even, there are six minimum-distance codewords (encoded bit strings) when the parity is odd. To construct level-2 codewords using the level-1 minimum-distance encoded bit strings, we first choose five level-1 blocks from the six ones of each level-2 block, and then pick one of the minimum-distance encoded bit strings from each of the five level-1 blocks. The encoded bit string of the other level-1 block is automatically determined by the parity-check condition corresponding to the $Z$-stabilizer. In general, the determined encoded bit string is not included in the minimum-distance ones of this block. So we evaluate the distance of this encoded bit string, and then obtain the distance of the level-2 block by summing the distances of the six level-1 blocks. We select the minimum-distance codewords (encoded bit strings) of the level-2 block among all the choices of the five level-1 blocks and their encoded bit strings, and keep only them as minimum-distance candidates. This is our strategy to utilize the minimum-distance candidates while satisfying the parity-check condition. Doing this selection recursively, we finally obtain the minimum-distance candidates at the logical level. If we have multiple candidates at the logical level, we randomly choose one of them. For error detection used in the proposed encoders, the decoder returns $F$ unless all the numbers of the candidates at the levels from $L_{D}$ to $L$ are one, where $L$ is the logical level and $L_{D}$ ($\leq L$) is chosen appropriately.
In the following, we explain our proposed decoding method in more detail using flowcharts. As an example, we mainly focus on the level-3 case. Figure 7 shows the whole process of the decoding in the level-3 case. The subroutines Sub1 and Sub2 are explained above. So in the following, we explain the subroutine Sub3 in detail.
Figure 8 shows the process of Sub3. In the subroutine Sub31, we first construct level-3 codewords by determining the encoded bit string of the $b_{2}$th level-2 block ($b_{2}=1,\ldots,6$) using the minimum-distance encoded bit strings of the other five level-2 blocks determined in Sub2 according to the parity-check condition ($Z$-stabilizer condition). In general, the level-2 encoded bit string determined here is not included in the minimum-distance encoded bit strings of this block determined in Sub2. So in the subroutine Sub32 explained later, we evaluate the distance of this encoded bit string. Then we evaluate the distances of the level-3 codewords by summing the distances of their six level-2 blocks. We finally select level-3 encoded bit strings corresponding to the level-3 codewords with minimum distance.
Figure 9 shows the details of Sub31 when $b_{2}=6$, where $N_{1}$ to $N_{5}$ denote the numbers of the minimum-distance candidates of the first to fifth level-2 blocks, respectively. If the total number of candidates, $N_{1}N_{2}N_{3}N_{4}N_{5}$, is larger than a preset threshold $N_{\mathrm{th}3}$, we reduce the number by randomly choosing one of the candidates from the level-2 block with the largest $N$, in order to keep the decoding time short. In this work, we set $N_{\mathrm{th}3}$ to $10^{5}$. Then we determine the encoded bit string of the sixth level-2 block using the $n_{1}$th to $n_{5}$th minimum-distance encoded bit strings of the other five level-2 blocks, respectively, according to the parity-check condition. More concretely, the encoded bit string of the sixth level-2 block is determined as follows:
$x_{6,j,l}^{(2)}=x_{1,j,l}^{(2)}+x_{2,j,l}^{(2)}+x_{3,j,l}^{(2)}+x_{4,j,l}^{(2)}+x_{5,j,l}^{(2)},$ (37)
where the summation is modulo 2 and $j=1,\ldots,4,l=1,\ldots,4$. The distance of the encoded bit string determined here is evaluated in the subroutine Sub32. We finally construct the level-3 codeword using the encoded bit strings of the six level-2 blocks and evaluate the level-3 distance by summing the level-2 distances.
Figure 10 shows the subroutine Sub32 when $b_{2}=6$, where we evaluate the distance of the sixth level-2 block determined in Sub31, as mentioned above. In Fig. 10, $M_{1}$ to $M_{6}$ denote the numbers of the minimum-distance candidates of the first to sixth level-1 blocks, respectively, in the sixth level-2 block. If $M_{1}+M_{2}+M_{3}+M_{4}+M_{5}+M_{6}$ is larger than a preset threshold $M_{\mathrm{th}2}$, we reduce the number by randomly choosing one of the candidates from the level-1 block with the largest $M$. In this work, we set $M_{\mathrm{th}2}$ to 6. Then in the subroutine Sub33, we evaluate the distance of the sixth level-2 block when the minimum-distance candidates of the $b_{1}$th level-1 block ($b_{1}=1,\ldots,6$) are used. We finally select the minimum distance among all the choices of the level-1 blocks and their minimum-distance candidates.
Figure 11 shows the subroutine Sub33 when $b_{1}=1$. We first determine the encoded bit strings of the five level-1 blocks other than the first level-1 block in the sixth level-2 block using the m1th minimum-distance encoding bit string of the first level-1 block determined in Sub1 and the level-2 encoding bit string determined in Sub31 according to the parity-check condition and the definition of the encoded $Z$ operator. More concretely, the encoded bit strings of the five level-1 blocks are determined as follows:
$x_{6,2,l}^{(1)}$ $=x_{6,1,l}^{(1)}+x_{6,1,l}^{(2)},$ (38)
$x_{6,3,l}^{(1)}$ $=x_{6,2,l}^{(1)}+x_{6,2,l}^{(2)},$ (39)
$x_{6,5,l}^{(1)}$ $=x_{6,2,l}^{(1)}+x_{6,1,l}^{(2)}+x_{6,2,l}^{(2)}+x_{6,3,l}^{(2)}+x_{6,4,l}^{(2)},$ (40)
$x_{6,4,l}^{(1)}$ $=x_{6,5,l}^{(1)}+x_{6,3,l}^{(2)},$ (41)
$x_{6,6,l}^{(1)}$ $=x_{6,5,l}^{(1)}+x_{6,4,l}^{(2)},$ (42)
where the summation is modulo 2 and $l=1,\ldots,4$. We obtain the distance of the sixth level-2 block by summing the distances of the six level-1 blocks.
The decoding of the level-4 C642 is done in a similar manner to the above level-3 case. In the level-4 case, we set the level-4 threshold $N_{\mathrm{th}4}$ corresponding to the above level-3 threshold $N_{\mathrm{th}3}$ to the same value $10^{5}$ and the level-3 threshold $M_{\mathrm{th}3}$ corresponding to the above level-2 threshold $M_{\mathrm{th}2}$ to 12.
Appendix G Simulation of the bit-flip error model
To evaluate the performance of the decoders, we did the following numerical Monte Carlo simulation. We first prepare an error-free logical all-zero state of the C642 using the encoder in Fig. 5d. Then independent physical-qubit bit-flip errors are induced with probability $p_{\mathrm{flip}}$ per physical qubit. Finally, we ideally measure the physical qubits in the $Z$ basis and decode the outcomes. If all the logical-qubit outcomes are 0, the decoding succeeds, and otherwise it fails. We evaluated the logical failure probabilities by sufficiently many trials of this simulation.
Appendix H Error-detection gadgets
Figures 12a and 12b show the level-1 $Z$-error and $X$-error error-detection gadgets, respectively, used in the level-2 zero-state encoder in Fig. 3b. Figures 12c and 12d show the level-2 $Z$-error and $X$-error error-detection gadgets, respectively, used in the level-3 zero-state encoder.
Appendix I Simulation of logical CNOT gates
To evaluate the performance of full transversal logical CNOT gates, we did the numerical Monte Carlo simulation shown in Fig. 13. From many trials of this simulation, we estimate the error probability and its standard error of 10 sets of full transversal logical CNOT gates, which are denoted by $p_{10}$ and $\Delta_{10}$, respectively. Then we evaluate those for one set, which are denoted by $p_{1}$ and $\Delta_{1}$, as $p_{1}=1-(1-p_{10})^{1/10}$ and $\Delta_{1}=\frac{\Delta_{10}}{10}(1-p_{10})^{1/10-1}$. Since each set includes $K^{(L)}$ logical CNOT gates at level $L$, we finally evaluate the error probability and its standard error per logical CNOT gate denoted by $p_{\mathrm{CNOT}}$ and $\Delta_{\mathrm{CNOT}}$ as $p_{\mathrm{CNOT}}=1-(1-p_{1})^{1/K^{(L)}}$ and $\Delta_{\mathrm{CNOT}}=\Delta_{1}/K^{(L)}(1-p_{1})^{1/K^{(L)}-1}$.
Appendix J Logical phase and non-Clifford gates
The full transversal logical phase and non-Clifford gates, $S$ and ${R_{Y}(\pi/4)}$, are performed with logical $|Y\rangle=(|0\rangle+i|1\rangle)/\sqrt{2}$ and $|H\rangle=\cos(\pi/8)|0\rangle+\sin(\pi/8)|0\rangle$, respectively, as shown in Figs. 14a and  14b. For the non-Clifford gate, we need partial transversal logical Hadamard gates depending on the measurement outcomes. We can also perform partial transversal logical phase and non-Clifford gates by replacing the full transversal logical Hadamard gates with partial ones, because then the two CNOT gates on the qubits without Hadamard gates are cancelled out.
Appendix K Partial transversal logical Hadamard gates
The partial transversal logical Hadamard gates are performed with a logical partial |+ket|$\rangle$| + ⟩ state and full transversal logical CNOT and Hadamard gates, as shown in Fig. 15a. The logical partial |+ket|$\rangle$| + ⟩ states are prepared fault-tolerantly by the arbitrary-state encoder in Fig. 5e and the standard state verification shown in Fig. 15b.
Appendix L Logical SWAP gates
The level-1 encoded SWAP gates are performed by physical SWAP gates as shown in Fig. 5c. At higher levels, e.g., level 3, simultaneous physical SWAP gates $\{\mathrm{Q}_{i^{\prime},j^{\prime},1}\leftrightarrow\mathrm{Q}_{i^{\prime},j^{\prime},3}|i^{\prime}=1,\ldots,6,j^{\prime}=1,\ldots,6\}$ result in simultaneous logical SWAP gates $\{\mathrm{Q}_{i,j,1}^{(3)}\leftrightarrow\mathrm{Q}_{i,j,2}^{(3)}|i=1,\ldots,4,j=1,\ldots,4\}$. Similarly, the other simultaneous logical SWAP gates, such as $\{\mathrm{Q}_{i,3,k}^{(3)}\leftrightarrow\mathrm{Q}_{i,4,k}^{(3)}|i=1,\ldots,4,k=1,\ldots,4\}$ and $\{(\mathrm{Q}_{1,j,k}^{(3)},\mathrm{Q}_{2,j,k}^{(3)})\leftrightarrow(\mathrm{Q}_{3,j,k}^{(3)},\mathrm{Q}_{4,j,k}^{(3)})|j=1,\ldots,4,k=1,\ldots,4\}$, are performed by corresponding simultaneous physical SWAP gates, such as $\{\mathrm{Q}_{i^{\prime},4,k^{\prime}}\leftrightarrow\mathrm{Q}_{i^{\prime},6,k^{\prime}}|i^{\prime}=1,\ldots,6,k^{\prime}=1,\ldots,6\}$ and $\{(\mathrm{Q}_{1,j^{\prime},k^{\prime}},\mathrm{Q}_{2,j^{\prime},k^{\prime}},\mathrm{Q}_{3,j^{\prime},k^{\prime}})\leftrightarrow(\mathrm{Q}_{4,j^{\prime},k^{\prime}},\mathrm{Q}_{5,j^{\prime},k^{\prime}},\mathrm{Q}_{6,j^{\prime},k^{\prime}})|j^{\prime}=1,\ldots,6,k^{\prime}=1,\ldots,6\}$. In general, we can perform simultaneous logical SWAP gates between the logical qubits corresponding to hypercubes on two parallel hyperplanes by simultaneous physical SWAP gates.
Appendix M Logical CNOT gates
Partial transversal logical CNOT gates between two registers (inter-block logical CNOT gates) can be performed using two ancilla registers in the same partial |+ket|$\rangle$| + ⟩ state and full transversal logical CNOT and Hadamard gates, as shown in Fig. 16a. Using logical SWAP gates, we can perform a logical CNOT gate on an arbitrary logical-qubit pair between the two registers. On the other hand, logical CNOT gates in a register (intra-block logical CNOT gates) can be performed with an additional ancilla register in the logical all-zero state, logical SWAP gates, and full and partial transversal logical CNOT gates, as shown in Fig. 16b.